differentiate the following function with respect to x
sin^-1(sqrt(1-x/2)),0

Grade:12

1 Answers

Rinkoo Gupta
9 years ago

y=sin-1v(1-x2)

dy/dx=1/v1-(1-x2) .{-2x/v(1-x2)}

=-1/v(1-x2)

Thanks & Regards

Rinkoo Gupta

AskIITians Faculty